# Vevo2: Bridging Controllable Speech and Singing Voice Generation via Unified Prosody Learning

### v1

@@ 4 Experiments
Table 1: Results on zero-shot TTS task. (PT?: Whether post-training is applied.)
@@
Ground Truth - - - 10.91 0.687 1.47 $_{\scriptscriptstyle\pm\text{0.24}}$ 0.87 $_{\scriptscriptstyle\pm\text{0.19}}$ 12.65 0.658 1.04 $_{\scriptscriptstyle\pm\text{0.18}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$
F5-TTS Chen et al. (2025) - 101K 11.77 0.695 -1.05 $_{\scriptscriptstyle\pm\text{0.18}} -$0.06 $_{\scriptscriptstyle\pm\text{0.17}}$ 16.12 0.597 -1.92 $_{\scriptscriptstyle\pm\text{0.06}} -$1.53 $_{\scriptscriptstyle\pm\text{0.08}}$
MaskGCT Wang et al. (2025a) 50 101K 13.42 0.736 -1.14 $_{\scriptscriptstyle\pm\text{0.21}}$ 0.07 $_{\scriptscriptstyle\pm\text{0.31}}$ 11.71 0.753 -1.83 $_{\scriptscriptstyle\pm\text{0.14}} -$1.14 $_{\scriptscriptstyle\pm\text{0.11}}$
CosyVoice 2 Du et al. (2024a) 25 167K 11.20 0.706 0.10 $_{\scriptscriptstyle\pm\text{0.29}} -$0.16 $_{\scriptscriptstyle\pm\text{0.13}}$ 16.18 0.659 -1.63 $_{\scriptscriptstyle\pm\text{0.17}} -$0.94 $_{\scriptscriptstyle\pm\text{0.13}}$
@@
101K 15.52 0.677 -0.84 $_{\scriptscriptstyle\pm\text{0.15}} -$0.06 $_{\scriptscriptstyle\pm\text{0.27}}$ 19.39 0.658 -1.15 $_{\scriptscriptstyle\pm\text{0.13}} -$0.65 $_{\scriptscriptstyle\pm\text{0.18}}$
7K 37.57 0.611 -1.25 $_{\scriptscriptstyle\pm\text{0.26}} -$0.57 $_{\scriptscriptstyle\pm\text{0.17}}$ 25.38 0.692 -0.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.23 $_{\scriptscriptstyle\pm\text{0.14}}$
101K, 7K 14.32 0.681 -0.49 $_{\scriptscriptstyle\pm\text{0.24}} -$0.03 $_{\scriptscriptstyle\pm\text{0.19}}$ 15.78 0.708 -0.66 $_{\scriptscriptstyle\pm\text{0.16}} -$0.17 $_{\scriptscriptstyle\pm\text{0.24}}$
@@
    #Hz: Frame rate of the content-style tokenizer. Data: Pre-training data, where blue and red denote speech and singing data. The results on regular speech domain are in Appendix C.1. The best and the second best result (excluding ground truth) is shown in bold and by underlined.
4.1 Experimental Setup
Training Data For pre-training, we use Emilia He et al. (2024) as the speech data (101K hours), and follow SingNet’s pipeline Gu et al. (2025) to prepare 7K hours of singing voice data (source-separated from in-the-wild songs from the Internet). For post-training, we utilize INTP Zhang et al. (2025b) (which contains 250K preference speech pairs) to train the intelligibility reward model. For the proposed multi-objective alignment task, we employ three types of prosody prompts: (1) speech data: 20K positive samples randomly selected from INTP, (2) singing voice data: 20K samples from M4Singer Zhang et al. (2022a), and (3) instrumental sounds: rendered from MIDIs in M4Singer, covering 16 instruments.
Evaluation Data We consider various settings to construct the evaluation set: (1) For regular speech, we utilize SeedTTS test sets Anastassiou et al. (2024a), which is widely adopted in the TTS field Wang et al. (2025a); Du et al. (2024b); Chen et al. (2025). (2) For expressive speech, we employ Genshin-Voice Simon3000 (2025), which consists of professional voice acting recordings from the game Genshin Impact. (3) For singing voice, we use GTSinger Zhang et al. (2024b) and SingStyle111 Dai et al. (2023), which encompass a wide range of singing techniques, styles, and timbre types. (4) Additionally, for two unique tasks: humming-to-singing and instrument-to-singing, we utilize humming Liu et al. (2024) and instrumental sound including flute Braga Brum (2018) and violin Dong et al. (2022) for the evaluation. All evaluations are conducted using target text in both English and Chinese languages.
[…]
4.2 Mutual Benefits of Unified Modeling for Speech and Singing Voice
@@
To investigate whether speech and singing voice generation can be mutually beneficial through unified modeling, we examine the impact of different pre-training data distributions on Vevo2. Specifically, we conduct experiments on zero-shot TTS tasks and evaluate the performance using reference prompts from different distributions. For comparison, we select models with varying content-style tokenizer frame rates: MaskGCT Wang et al. (2025a) (50 Hz), CosyVoice 2 Du et al. (2024b) (25 Hz), and F5-TTS Chen et al. (2025) (a single-stage model without content-style tokenizer). The results are presented in Table 1.
@@
The experimental results demonstrate the following findings: (1) Comparing Vevo2 models with different pre-training data, we observe that on expressive speech domain, jointly using speech and singing data for pre-training improves all metrics (WER: 15.52 $\rightarrow$ 14.32, SIM: 0.677 $\rightarrow$ 0.681, N-CMOS: -0.84 $\rightarrow$ -0.49, SS-CMOS: -0.06 $\rightarrow$ -0.03). On singing voice domain, we observe that using only singing voice data for pre-training yields lower intelligibility compared to using speech data alone (WER: 25.38 vs. 19.39), likely due to the smaller size of the singing voice corpus. Furthermore, when combining speech and singing voice data for pre-training, Vevo2 demonstrates superior intelligibility (WER: 19.39 $\rightarrow$ 15.78) and improved naturalness and similarity (N-CMOS: -1.15 $\rightarrow$ -0.66, SS-CMOS: -0.65 $\rightarrow$ -0.17). These findings provide strong evidence that speech and singing voice generation can be mutually beneficial through unified pre-training. (2) Regarding the effect of post-training on Vevo2, our proposed approach brings further overall improvements to the pre-trained model, including enhanced intelligibility, speaker similarity, and naturalness. This demonstrates the effectiveness and significant potential of post-training techniques in controllable speech and singing voice generation. (3) Compared to existing baselines, the post-trained Vevo2 matches and even outperforms them on certain metrics in the expressive speech domain, despite its lower frame rate in the content-style tokenizer. More significantly, on singing voice domain, Vevo2 substantially outperforms existing zero-shot TTS models in subjective evaluations (N-CMOS: gaps greater than 1.5; SS-CMOS: gaps greater than 0.9), benefiting greatly from the incorporation of singing voice data during training. This validates Vevo2’s effectiveness as a unified controllable speech and singing voice generation model.
4.3 Controllability of Vevo2
@@
Table 2: Results on zero-shot SVS, speech editing, and singing lyric editing tasks.
Zero-Shot Singing Voice Synthesis
@@
Ground Truth 17.89 0.640 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 7.40 0.675 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
StyleSinger Zhang et al. (2024d) - - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 14.04 0.555 -0.37 $_{\scriptscriptstyle\pm\text{0.16}}$ 2.33 $_{\scriptscriptstyle\pm\text{0.36}}$
TCSinger Zhang et al. (2024c) 40.83 0.605 -0.09 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.55 $_{\scriptscriptstyle\pm\text{0.27}}$ 14.26 0.651 -0.28 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.58 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 60.83 0.586 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 23.60 0.653 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 24.77 0.643 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.29 $_{\scriptscriptstyle\pm\text{0.26}}$ 9.83 0.669 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.44 $_{\scriptscriptstyle\pm\text{0.18}}$
Zero-Shot Editing
@@
WER SIM N-CMOS PS-CMOS WER SIM N-CMOS MS-CMOS
SSR-Speech Wang et al. (2025b) 29.56 0.588 -1.36 $_{\scriptscriptstyle\pm\text{0.09}} -$0.89 $_{\scriptscriptstyle\pm\text{0.14}}$ 51.92 0.710 -1.43 $_{\scriptscriptstyle\pm\text{0.17}} -$1.10 $_{\scriptscriptstyle\pm\text{0.08}}$
F5-TTS Chen et al. (2025) 21.35 0.733 -0.19 $_{\scriptscriptstyle\pm\text{0.12}} -$0.21 $_{\scriptscriptstyle\pm\text{0.16}}$ 29.78 0.784 -1.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.97 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 23.54 0.795 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 29.69 0.841 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 16.83 0.799 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 17.98 0.848 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
@@
    Vevo2-base is pre-trained on both speech and singing data, corresponding to Vevo2-base (101K, 7K) in Table 1.
@@
Text and Prosody Control To evaluate Vevo2’s controllability of text and prosody (melody), we conduct various tasks including zero-shot SVS, speech editing, singing lyric editing, humming-to-singing, and instrument-to-singing (the latter two tasks will be discussed in Section 4.4).
@@
In the context of zero-shot SVS, given a target text, a music notation like MIDI, and a reference singer’s waveform, the model is required to generate singing voice that matches the content of the target text and music notation while preserving the timbre and style of the reference singer Zhang et al. (2024d, c); Zhao et al. (2025). For Vevo2, we render MIDI into piano waveform and use it as the prosodic source, which can also be considered as instrument-to-sing. Note that we also conduct instrument-to-sing experiments using real instrumental sound recordings, which will be discussed in Section 4.4. The results are shown in Table 2, where we include StyleSinger Zhang et al. (2024d) and TCSinger Zhang et al. (2024c) as two baselines. It demonstrates that: (1) After post-training, Vevo2 shows improved text-following ability and speaker similarity in both English and Chinese. We suspect that this is because that during pre-training, Vevo2 had not encountered prosody tokens extracted from instrumental sounds, resulting in unstable performance. By post-training, we not only supplement with MIDI-rendered instrumental sound data, but also adopt the specific alignment learning, effectively eliciting the model’s potential. (2) Compared to the baselines, the post-trained Vevo2 demonstrates superior text-following ability (WER) and naturalness (N-CMOS). Regarding melody-following, although the post-trained Vevo2’s Melody-MOS is slightly lower than TCSinger’s, its scores for both English and Chinese exceed 2.0, indicating that Vevo2 is at least capable of “roughly following the melody contour” (as defined by Melody-CMOS score of 2.0) in most cases.
@@
In addition to SVS, we adopt editing tasks to evaluate Vevo2’s text and prosody controllability. Vevo2’s advantage in editing tasks stems from its use of prosody tokens as explicit control, enabling it to preserve the prosody (melody) of the raw voice during editing. For instance, in singing lyric editing tasks Zhang et al. (2022c); Hong et al. (2023), Vevo2 can modify only the lyrics while maintaining the original singing voice melody. The results are shown in Table 2, where we include two speech editing baselines, SSR-Speech Wang et al. (2025b) and F5-TTS Chen et al. (2025). It can be observed that compared to the baselines, the post-trained Vevo2 not only achieves better intelligibility (WER), naturalness (N-CMOS), and speaker similarity (SIM), but also demonstrates superior prosody preservation of the original voice (PS-CMOS), particularly in the singing voice domain (MS-CMOS). We will provide more audio samples at our website to further demonstrate Vevo2’s advantages in singing lyric editing tasks.
@@
Table 3: Results on zero-shot SVC task.
@@
FACodec Ju et al. (2024) 32.81 0.434 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 36.97 0.473 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo-FM Zhang et al. (2025a) 24.05 0.567 -0.64 $_{\scriptscriptstyle\pm\text{0.07}} -$0.45 $_{\scriptscriptstyle\pm\text{0.16}}$ 0.71 $_{\scriptscriptstyle\pm\text{0.21}}$ 22.85 0.610 -0.55 $_{\scriptscriptstyle\pm\text{0.19}} -$0.44 $_{\scriptscriptstyle\pm\text{0.08}}$ 0.88 $_{\scriptscriptstyle\pm\text{0.27}}$
CosyVoice2-FM Du et al. (2024b) 23.59 0.553 -0.64 $_{\scriptscriptstyle\pm\text{0.11}} -$0.63 $_{\scriptscriptstyle\pm\text{0.18}}$ 2.25 $_{\scriptscriptstyle\pm\text{0.15}}$ 20.14 0.589 -0.63 $_{\scriptscriptstyle\pm\text{0.15}} -$0.54 $_{\scriptscriptstyle\pm\text{0.07}}$ 1.65 $_{\scriptscriptstyle\pm\text{0.23}}$
NeuCoSVC 2 Sha et al. (2024) 30.61 0.481 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 32.26 0.519 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
SeedVC Liu (2024) 22.86 0.508 -0.05 $_{\scriptscriptstyle\pm\text{0.13}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$ 2.89 $_{\scriptscriptstyle\pm\text{0.09}}$ 15.65 0.550 -0.33 $_{\scriptscriptstyle\pm\text{0.21}} -$0.54 $_{\scriptscriptstyle\pm\text{0.14}}$ 2.93 $_{\scriptscriptstyle\pm\text{0.14}}$
@@
    Vevo2-FM denotes the flow-matching model of Vevo2. The best and the second best result (excluding ground truth) is shown in bold and by underlined.

### v2

@@ V Results and Analysis
TABLE I: Results on zero-shot TTS task. (PT?: Whether post-training is applied.)
@@
Ground Truth - - - 10.91 0.687 1.47 $_{\scriptscriptstyle\pm\text{0.24}}$ 0.87 $_{\scriptscriptstyle\pm\text{0.19}}$ 12.65 0.658 1.27 $_{\scriptscriptstyle\pm\text{0.14}}$ 0.55 $_{\scriptscriptstyle\pm\text{0.16}}$
F5-TTS [f5tts] - 101K 11.77 0.695 -1.05 $_{\scriptscriptstyle\pm\text{0.18}} -$0.06 $_{\scriptscriptstyle\pm\text{0.17}}$ 16.12 0.597 -1.89 $_{\scriptscriptstyle\pm\text{0.07}} -$1.29 $_{\scriptscriptstyle\pm\text{0.10}}$
MaskGCT [maskgct] 50 101K 13.42 0.736 -1.14 $_{\scriptscriptstyle\pm\text{0.21}}$ 0.07 $_{\scriptscriptstyle\pm\text{0.31}}$ 11.71 0.753 -1.74 $_{\scriptscriptstyle\pm\text{0.14}} -$0.97 $_{\scriptscriptstyle\pm\text{0.13}}$
CosyVoice 2 [cosyvoice] 25 167K 11.20 0.706 0.10 $_{\scriptscriptstyle\pm\text{0.29}} -$0.16 $_{\scriptscriptstyle\pm\text{0.13}}$ 16.18 0.659 -1.68 $_{\scriptscriptstyle\pm\text{0.15}} -$1.12 $_{\scriptscriptstyle\pm\text{0.16}}$
@@
101K 15.52 0.677 -0.84 $_{\scriptscriptstyle\pm\text{0.15}} -$0.06 $_{\scriptscriptstyle\pm\text{0.27}}$ 19.39 0.658 -1.02 $_{\scriptscriptstyle\pm\text{0.16}} -$0.48 $_{\scriptscriptstyle\pm\text{0.18}}$
7K 37.57 0.611 -1.25 $_{\scriptscriptstyle\pm\text{0.26}} -$0.57 $_{\scriptscriptstyle\pm\text{0.17}}$ 25.38 0.692 -0.69 $_{\scriptscriptstyle\pm\text{0.26}} -$0.36 $_{\scriptscriptstyle\pm\text{0.17}}$
101K, 7K 14.32 0.681 -0.49 $_{\scriptscriptstyle\pm\text{0.24}} -$0.03 $_{\scriptscriptstyle\pm\text{0.19}}$ 15.78 0.708 -0.28 $_{\scriptscriptstyle\pm\text{0.14}} -$0.09 $_{\scriptscriptstyle\pm\text{0.21}}$
@@
    #Hz: Frame rate of the content-style tokenizer. Data: Pre-training data, where blue and red denote speech and singing data. The results on the regular speech domain are provided in the supplementary materials. The best and the second best result (excluding ground truth) are shown in bold and by underlined.
V-A Benefits of Unified Modeling for Speech and Singing Voice
@@
To investigate whether speech and singing voice generation can be mutually beneficial through unified modeling, we examine the impact of different pre-training data distributions on Vevo2. Specifically, we conduct experiments on zero-shot TTS tasks and evaluate the performance using reference prompts from different distributions. For comparison, we select models with varying content-style tokenizer frame rates: MaskGCT [maskgct] (50 Hz), CosyVoice 2 [cosyvoice2] (25 Hz), and F5-TTS [f5tts] (a single-stage model without content-style tokenizer). The results are presented in Table I, which demonstrates the following findings:
@@
(1) Speech and singing voice generation mutually benefit from unified pre-training: Comparing Vevo2 models with different pre-training data, we observe that on expressive speech domain, jointly using speech and singing data for pre-training improves all metrics (WER: 15.52 $\rightarrow$ 14.32, SIM: 0.677 $\rightarrow$ 0.681, N-CMOS: -0.84 $\rightarrow$ -0.49, SS-CMOS: -0.06 $\rightarrow$ -0.03). On singing voice domain, we observe that using only singing voice data for pre-training yields lower intelligibility compared to using speech data alone (WER: 25.38 vs. 19.39), likely due to the smaller size of the singing voice corpus. Furthermore, when combining speech and singing voice data for pre-training, Vevo2 demonstrates superior intelligibility (WER: 19.39 $\rightarrow$ 15.78) and improved naturalness and similarity (N-CMOS: -1.02 $\rightarrow$ -0.28, SS-CMOS: -0.48 $\rightarrow$ -0.09).
(2) Effectiveness of the proposed post-training: The proposed post-training brings further overall improvements to the pre-trained model (Vevo2-base), including enhanced intelligibility (WER), speaker similarity (SIM and SS-CMOS), and naturalness (N-CMOS). This demonstrates the effectiveness and significant potential of post-training techniques in controllable speech and singing voice generation.
(3) Competitive performance of Vevo2 compared to existing models: Compared to the baselines, the post-trained Vevo2 matches and even outperforms them on certain metrics in the expressive speech domain, despite its lower frame rate in the content-style tokenizer. More significantly, on singing voice domain, Vevo2 substantially outperforms existing zero-shot TTS models in subjective evaluations (N-CMOS: gaps greater than 1.5; SS-CMOS: gaps greater than 0.9), benefiting greatly from the incorporation of singing voice data during training. This validates Vevo2’s effectiveness as a unified controllable speech and singing voice generation model.
V-B Controllability over Text and Prosody
@@
TABLE II: Results on zero-shot SVS, speech editing, and singing lyric editing tasks.
Singing Voice Synthesis
@@
Ground Truth 17.89 0.640 0.815 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 7.40 0.675 0.782 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
StyleSinger [stylesinger] - - - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 14.04 0.555 0.670 -0.37 $_{\scriptscriptstyle\pm\text{0.16}}$ 2.33 $_{\scriptscriptstyle\pm\text{0.36}}$
TCSinger [tcsinger] 40.83 0.605 0.724 -0.09 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.55 $_{\scriptscriptstyle\pm\text{0.27}}$ 14.26 0.651 0.707 -0.28 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.58 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 60.83 0.586 0.719 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 23.60 0.653 0.674 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 24.77 0.643 0.752 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.29 $_{\scriptscriptstyle\pm\text{0.26}}$ 9.83 0.669 0.710 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.44 $_{\scriptscriptstyle\pm\text{0.18}}$
Editing
@@
WER SIM FPC N-CMOS PS-CMOS WER SIM FPC N-CMOS MS-CMOS
SSR-Speech [ssrspeech] 29.56 0.588 0.721 -1.36 $_{\scriptscriptstyle\pm\text{0.09}} -$0.89 $_{\scriptscriptstyle\pm\text{0.14}}$ 51.92 0.710 0.804 -1.43 $_{\scriptscriptstyle\pm\text{0.17}} -$1.10 $_{\scriptscriptstyle\pm\text{0.08}}$
F5-TTS [f5tts] 21.35 0.733 0.730 -0.19 $_{\scriptscriptstyle\pm\text{0.12}} -$0.21 $_{\scriptscriptstyle\pm\text{0.16}}$ 29.78 0.784 0.821 -1.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.97 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 23.54 0.795 0.782 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 29.69 0.841 0.872 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 16.83 0.799 0.792 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 17.98 0.848 0.877 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
@@
    Vevo2-base here is pre-trained on both speech and singing data, corresponding to the row “Vevo2-base (101K, 7K)" in Table I.
@@
To evaluate Vevo2’s controllability with respect to text and prosody (melody), we conduct various tasks including zero-shot SVS, speech editing, singing lyric editing, humming-to-singing, and instrument-to-singing (the latter two tasks will be discussed in Section V-F).
Zero-Shot SVS Given a target text, a music notation like MIDI, and a reference singer’s waveform, a zero-shot SVS model is required to generate singing voice that matches the content of the target text and music notation while preserving the timbre and style of the reference singer [stylesinger, tcsinger, spsinger]. For Vevo2, we render MIDI into piano waveform and use it as the prosodic source, which can also be considered as instrument-to-singing. The results are shown in Table II, where we include StyleSinger [stylesinger] and TCSinger [tcsinger] as two baselines.
@@
It demonstrates that: (1) After post-training, Vevo2 shows improved text-following ability and speaker similarity in both English and Chinese. We suspect that this is because that during pre-training, Vevo2 had not encountered prosody tokens extracted from instrumental sounds, resulting in unstable performance. By post-training, we not only supplement with MIDI-rendered instrumental sound data, but also adopt the specific alignment learning, effectively eliciting the model’s potential. (2) Compared to the baselines, the post-trained Vevo2 demonstrates superior text-following ability (WER) and naturalness (N-CMOS). Regarding melody-following, although the post-trained Vevo2’s Melody-MOS is slightly lower than TCSinger’s, its scores for both English and Chinese exceed 2.0, indicating that Vevo2 is at least capable of “roughly following the melody contour” (as defined by Melody-CMOS score of 2.0) in most cases.
Zero-Shot Editing In addition to SVS, we adopt editing tasks to evaluate Vevo2’s text and prosody controllability. The used evaluation sets are constructed with DeepSeek-V3 [deepseek-v3]. Specifically, we utilized the same expressive speech and singing voice evaluation data from our zero-shot TTS tasks. For each sample’s text, we generated edited versions using DeepSeek-V3 with the following prompt: "Please edit the following text by adding, deleting, or modifying content as needed. Note: (1) Ignore any existing punctuation errors in the original text and ensure proper punctuation in the edited version. (2) The edited version must differ from the original text. (3) Return only the final edited text without any additional explanations."
@@
Vevo2’s advantage in editing tasks stems from its use of prosody tokens as explicit control, enabling it to preserve the prosody (melody) of the raw voice during editing. For instance, in singing lyric editing tasks [editsinger, unisinger], Vevo2 can modify only the lyrics while maintaining the original singing voice melody. The results are shown in Table II, where we include two speech editing baselines, SSR-Speech [ssrspeech] and F5-TTS [f5tts]. It can be observed that compared to the baselines, the post-trained Vevo2 not only achieves better intelligibility (WER), naturalness (N-CMOS), and speaker similarity (SIM), but also demonstrates superior prosody preservation of the original voice (PS-CMOS), particularly in the singing voice domain (MS-CMOS). We will provide more audio samples at our website to further demonstrate Vevo2’s advantages in singing lyric editing tasks.
@@
TABLE III: Results on zero-shot SVC task. Vevo2-FM denotes the flow-matching model of Vevo2.
@@
FACodec [naturalspeech3] 32.81 0.434 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 36.97 0.473 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo-FM [vevo] 24.05 0.567 -0.64 $_{\scriptscriptstyle\pm\text{0.07}} -$0.45 $_{\scriptscriptstyle\pm\text{0.16}}$ 0.71 $_{\scriptscriptstyle\pm\text{0.21}}$ 22.85 0.610 -0.55 $_{\scriptscriptstyle\pm\text{0.19}} -$0.44 $_{\scriptscriptstyle\pm\text{0.08}}$ 0.88 $_{\scriptscriptstyle\pm\text{0.27}}$
CosyVoice2-FM [cosyvoice2] 23.59 0.553 -0.64 $_{\scriptscriptstyle\pm\text{0.11}} -$0.63 $_{\scriptscriptstyle\pm\text{0.18}}$ 2.25 $_{\scriptscriptstyle\pm\text{0.15}}$ 20.14 0.589 -0.63 $_{\scriptscriptstyle\pm\text{0.15}} -$0.54 $_{\scriptscriptstyle\pm\text{0.07}}$ 1.65 $_{\scriptscriptstyle\pm\text{0.23}}$
NeuCoSVC 2 [neucosvc] 30.61 0.481 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 32.26 0.519 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
SeedVC (SVC) [seedvc] 22.86 0.508 -0.05 $_{\scriptscriptstyle\pm\text{0.13}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$ 2.89 $_{\scriptscriptstyle\pm\text{0.09}}$ 15.65 0.550 -0.33 $_{\scriptscriptstyle\pm\text{0.21}} -$0.54 $_{\scriptscriptstyle\pm\text{0.14}}$ 2.93 $_{\scriptscriptstyle\pm\text{0.14}}$
@@
    The best and the second best result (excluding ground truth) is shown in bold and by underlined.
TABLE IV: Results on zero-shot accent and emotion conversion tasks.
Model Accent Conversion Emotion Conversion
WER SIM A-SIM UTMOS WER SIM E-SIM UTMOS
Vevo [vevo] 30.37 0.702 0.526 3.49 13.31 0.697 0.406 3.95
Vevo2-base 10.08 0.760 0.537 3.72 11.36 0.700 0.400 3.82
Vevo2 7.73 0.768 0.529 3.76 7.89 0.714 0.402 3.89
    A-SIM: Accent-SIM. E-SIM: Emotion-SIM.
TABLE V: Effect of duration control of Vevo2.
Task DDUR ($\downarrow$)
DC ($\uparrow$)
SE 0.22s 97.1%
SLE 0.20s 98.4%
SVC 0.12s 98.4%
    SE: Speech Editing. SLE: Singing Lyric Editing.
V-C Controllability over Timbre and Style